SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 26 March 2005

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Saturday, March 26, 2005

image

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below. Saturday -- but not as big a crew rest day as originally planned...

Service Module Thermal Control System, Mission Operations Directorate, 20 Sep 2000 [Acrobat] This 53 page document covers operations and design features of the Service Module's Thermal System.

In a late change to their schedule, Sharipov and Chiao today replaced the #1 pump panel (4SPN1) of the Service Module (SM)'s internal cooling loop 2 (KOV-2). The panel failed early yesterday morning, causing a switchover to pump panel #2. Rather than isolating the failed pump, the entire 4SPN1 panel was replaced, to mitigate any contingencies from the thermal loop during EVA-13. [Each of the two internal KOV loops has two 4SPN panels with two pumps each, operating in series. It was not clear at the time if both pumps on 4SPN1 had failed or only one. Ever since the installation of a new design by Gennady Padalka on 7/16/04, both pumps in a panel can be replaced individually, whereas the old panels had the pumps hard-welded to them.]

In support for EVA-13 on Monday, CDR Chiao initiated the charging process on the NiMH (nickel metal hydride) batteries for the Orlan helmet lights in the BSA (battery stowage assembly) in the Joint Airlock. [Leroy will terminate battery charging tomorrow and then install the NiMHs in the helmet light assemblies. During the spacewalk, the lights will be left powered on from beginning to end, to allow their deep discharge in support of a 6-month maintenance requirement.]

Leroy completed the regular weekly maintenance reboot on the operational PCS (portable computer system) laptops, changing out a ROM on one, and the bi-monthly restart of the OCA comm router laptop.

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The CDR also collected the periodic reading of the cabin air's current CO2 partial pressure in the SM and Lab, using the new U.S. CDMK (carbon dioxide monitoring kit) #1013 (new backup is #1014) for calldown, along with the battery status, for use in trending analyses.

At ~9:55am EST, the crew held their weekly teleconference with ISS Program Management at JSC/Houston via S-band/audio.

Leroy and Salizhan completed their regular daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser, and VELO cycle with bungee cord load trainer, plus a session on CEVIS (cycle ergometer with vibration isolation) for Leroy. [Salizhan's daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO (today: Day 2 of a new set).]

Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, which he then erased on the HRM storage medium.

In preparation for the early-morning EVA-13 on Monday, the crew's sleep cycle will shift the day before. [After wakeup tomorrow morning (3/27) at the regular 1:00am EST, sleep time begins already 7 hours later (8:00am), followed by second wakeup on that day at 4:30pm. After spacewalk conclusion, bedtime on Monday starts after noon (12:30pm) and ends at 1:00am Tuesday, returning to the regular cycle.]

Weekly Science Update (Expedition Ten -- 21st):

GASMAP: Nothing new.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): Nothing new.

Hand Posture Analyzer (HPA: Nothing new.

In-Space Soldering Investigation (ISSI): Operations are complete. -->For a descriptive article on ISSI background & early surprising results, see AIAA's "Aerospace America" Magazine, March 2005 issue (page 24)

Pore Formation and Mobility Investigation (PFMI): Nothing new.

Space Acceleration Measurement System (SAMS): SAMS remains in nominal operations.

Microgravity Acceleration Measurement System (MAMS): MAMS remains in nominal operations.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Nothing new.

Binary Colloidal Alloy Test-3 (BCAT-3): Nothing new.

Renal Stone (RS): Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): In progress.

Dust and Aerosol Measurement Feasibility Test (DAFT): Nothing new.

Biotechnology Specimen Temperature Controller (BSTC): Nothing new.

Yeast Group Activation Packs (Yeast GAP): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Nothing new.

Earth Science Toward Exploration Research (ESTER): Nothing new.

Serial Network Flow Monitor (SNFM): Nothing new.

Fluid Merging Viscosity Measurement (FMVM): Nothing new.

Space Experiment Module (SEM): Nothing new.

Viscous Liquid Foam--Bulk Metallic Glass (Foam): Nothing new.

Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY): Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.

Miscible Fluids in Microgravity (MFMG): Nothing new.

Educational Payload Operations (EPO): Nothing new.

Capillary Flow Experiment (CFE): Nothing new.

Crew Earth Observations (CEO): As of 3/23, total number of CEO images downlinked is 13,286 frames. Recent excellent 800mm views of Lima, Peru are the best ever seen by experimenters. Weather, illumination, and the crew's fine technique combined to exceed requirements for this challenging target.

Today's optional CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Sobat fans, SE Sudan (several poorly described megafans have developed at the foot of the Ethiopian plateau. General views are still required to compare with other imagery. A series of overlapping images looking right of track at the base of the mountains was requested), Navassa Island reef, Caribbean (Nadir pass. Coral reef distribution [for mapping] and color [an indicator of coral health] are points of interest), Bahamas internal waves (looking right toward the sunglint point for any internal waves, especially between islands), and Yucatan smoke (Dynamic event. Although smoke from the present central American burning event is dissipating, source fires may still be visible [as fires per se, or as individual smoke plumes] right of track in Guatemala and the Yucatan peninsula of Mexico).

CEO photography can be viewed and studied at the websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

Upcoming Key Events (Eastern dates/times added):

  • EVA-13 -- 3/28 (hatch open~1:28am; hatch close~7:12am);
  • Soyuz TMA-6 (10S) launch -- 4/14 (8:46pm EDT); with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips & VC8 cosmonaut Roberto Vittori/ESA-Italy); launch time at Baikonur: 6:46am on 4/15.
  • Soyuz TMA-6 docking -- 4/16 (10:17pm EDT);
  • Soyuz TMA-5 (9S) undocking -- 4/24 (2:36pm EDT) with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS) and VC8 cosmonaut Vittori;
  • Soyuz TMA-5 landing -- 4/24 (6:01pm EDT (Kustanai: 4:01am on 4/25);.
  • LF1 (STS-114) launch -- 5/15;
  • Progress M-53 (18P) launch -- 6/10;
  • ULF1.1 (STS-121) launch -- NET 7/12;
  • Progress M-54 (19P) launch -- 8/24;
  • Soyuz TMA-7 (11S) launch -- 9/27.


ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.